# {cutout}04.2cm55cm1 empty TextMonkey     : An OCR-Free Large Multimodal Model for Understanding Document

### 07 Mar 2024 v1

@@ 4.2 Results
We conduct a comparative analysis of our approach with recent large multimodal models. For our evaluation, we utilize three Scene Text-Centric VQA datasets: STVQA [49], TextVQA [30], and OCRVQA [50]; three Document-Oriented VQA datasets: DocVQA [33], InfoVQA [51], and ChartQA [34]; and three Key Information Extraction (KIE) datasets: FUNSD [52], SROIE [53], and POIE [54]. To provide a thorough evaluation of performance, we also incorporate OCRBench [12] to produce an overall score.
Figure 5: Visualization results of TextMonkey. The bounding boxes generated by the model are visualized in red and integrated into the model’s input. The location of the ground truth are highlighted in green boxes for easy reference.
@@
Table V: Quantitative accuracy of text spotting. The “Total-Text” and “CTW1500” datasets do not use a specific vocabulary for evaluation, while the “ICDAR 2015” dataset uses a general vocabulary for evaluation of other models. TextMonkey is not fine-tuned by the downstream text spotting datasets without any vocabulary.
@@
TextMonkey 78.2 57.8 63.2 54.1 66.9 41.6
To further compare and assess the capabilities of our method, we conduct tests on additional datasets utilizing the specific evaluation metric provided in their paper: F1-score for Deepform and KLC, accuracy for WTQ, relaxed accuracy measure for ChartQA, and ANLS for DocVQA. The results, shown in Tab. IV, indicate that our model leads in performance on these datasets, outperforming other models, showing the model’s capability in handling documents, tables, and charts.
To show the extensive capabilities of our model, we assessed its performance on text spotting datasets without finetuning, as detailed in Tab. V. Given our model’s focus on identifying complete text passages, we segmented the predicted content into individual words for analysis. We employed two evaluation methodologies to evaluate our model’s performance. In the “Trans” mode, text is considered correct if the answer contains this word. Conversely, the “Pos” mode requires the consideration of positional information in accordance with the Mango [61]. For both metrics, due to granularity issues of the output (TextMonkey often produces an integrated paragraph while others only produce desired words), the metric can not strictly follow the evaluation setup; however, both should be quite similar, as both the error and correct situations match in calculations.
In terms of the “Trans” metric, our model surpassed SPTS v2 by a notable margin of 10.8%, demonstrating its superior understanding of textual content. In terms of the “Pos” metric, our model exhibited proficiency in zero-shot text spotting, underscoring its ability to interpret both the content and spatial positioning of text.
@@
4.3 Visualization
We conduct a qualitative evaluation of TextMonkey perception and comprehension capabilities across various scenarios, including natural scenes and document images. As shown in the left part of Fig. 5, TextMonkey accurately locates and identifies text in both scene and document images. Besides, natural images in Fig. 5 (a), documents in Fig. 5 (b), charts in Fig. 5 (c), and tables in Fig. 5 (d) exemplify TextMonkey’s adeptness at discerning and interpreting visual and textual information within a wide range of scenarios.
@@ 4.4 Ablation Study
Ablation study on strategies of reducing token length. As demonstrated in Tab. VIII, substituting important tokens with random ones (without token filter) leads to an average decline in performance by roughly 12.7%. This decline is attributed to the increased complexity of optimizing random queries, which necessitates more datasets to achieve a generalized representation compared to utilizing significant tokens. Solely focusing on pivotal features (without resampler) and directly eliminating features incurs a loss of some information, showing a decrease in performance, such as a 2.1% drop in SROIE. Additionally, neglecting the order of tokens (with unsorted token filter) does not markedly impair performance, owing to the language model’s inherent ability to organize unordered tokens. Nevertheless, the lack of token order can still lead to losses, especially evident in DocVQA outcomes, with a 2.2% decrease in performance.
Interaction between input resolution and the number of tokens remained. As shown in Tab. IX, Directly increasing the resolution without compressing tokens can actually lead to consistent performance loss, especially with a decrease of 9.2% performance in DocVQA. We speculate that the increase in resolution results in a significant increase in redundant tokens, making it more difficult to find crucial information in our setting. Therefore, compressing tokens reasonably can lead to higher performance. Considering the sparsity of information in large-sized images, it is also necessary to consider selecting an appropriate value of “r” for different input resolutions. Besides, increasing the input resolution brings benefits to the dataset, which contains many large-sized images, with 0.2% performance gain for DocVQA and 3.2% performance gain for InfoVQA. However, for datasets like TextVQA and SROIE, which contain much smaller images, increasing the input resolution directly does not yield any gains.
Table IX: Interaction between resolution and the number of tokens remained “r”.
@@
Figure 6: Examples of structuralization of chart and table.
@@ 4.5 Structuralization
The structuralization of charts and tables holds substantial practical value. Structured charts and tables present data in a clear format, and by extracting structural information from images, computers can accurately parse and extract the data. This makes data analysis, statistics, and modeling more efficient and precise. It also helps reduce the complexity of information and improves its comprehensibility. As depicted in Fig. 6, our model is capable of structuring charts and tables into JSON format, demonstrating its potential for downstream applications. According to Tab. IV, TextMonkey† exhibits a performance improvement of 1.3% and 1.4% on tables and charts, respectively. This underscores that high-quality data not only enables the model’s structuralization capabilities but also amplifies the effectiveness of the benchmarks.
@@ 4.6 App Agent
Recently, there has been a lot of attention on using LMMs for the task of acting as agents for smartphone applications  [62, 63]. Unlike existing intelligent phone assistants like Siri, which operate through system back-end access and function calls, this agent interacts with smartphone applications in a human-like manner, using low-level operations such as clicking and swiping on the graphical user interface (GUI). It eliminates the need for system back-end access, enhancing security and privacy as the agent does not require deep system integration. The GUI primarily consists of icons and text, and we explore the feasibility of TextMonkey on this aspect. We transformed 15k user click data from the Rico [64] dataset and performed downstream fine-tuning using TextMonkey. As qualitatively shown in Fig. 8, our model is able to understand user intent and click on the corresponding icons, which suggests the potential of the model to serve as an app agent by using downstream data.
Figure 7: Examples of grounding the position of the answer.
@@
4.7 Discussion
Interpretability. By examining the grounding information, we can identify the reasons behind the model’s errors, enhancing a better understanding of the model’s behavior. As shown in Fig. 7 (a), we ground to the white region, indicating that the model might be engaging in hallucination. We correctly identify the location but recognize it wrongly in Fig. 7 (b). In Fig. 7 (c), we ground to incorrect text but still provided the correct answer, suggesting that there might be randomness in the answers at this point. In Fig. 7 (d), the alignment between the position and text indicates that the model is more confident in its predictions. Therefore, based on these analyses, we can gain a better understanding of the model’s behavior and have a better awareness of the model’s hallucination, thus reducing the model’s hallucination.
Chain-of-Thought. We also conduct experiments on several datasets and observe inconsistent improvements if we require a model to provide the answer’s position, as shown in Tab. X. Upon examining the wrong answer, we find that the requirement of grounding might have partially affected certain reasoning needs. Furthermore, automating the process of constructing a thinking chain in subsequent steps could be a potential avenue for future work.
@@ 5 Conclusion
This paper introduces TextMonkey to address the challenges associated with text-heavy tasks such as document question answering and fine-grained text analysis. We adopt Shifted Window Attention with zero initialization to help establish relationships while increasing input resolutions using a sliding window. Increasing the resolution simultaneously increases the number of tokens. Through analyzing the redundancy of tokens, our proposed Token Resampler effectively reduces the number of tokens. Furthermore, by engaging in multiple text-oriented tasks simultaneously, TextMonkey enhances its perception and understanding of spatial relationships, leading to improved interpretability, reduced model hallucinations, and support for clicking screen-shots. By comparing our model with various LMMs, our model achieved excellent results on multiple benchmarks. It is worth mentioning that we also find that directly increasing the input resolution does not always lead to improvements, particularly for much smaller images. This underscores the necessity of creating an efficient method for scaling resolution in documents where size changes can be dramatic.

### 15 Mar 2024 v2

@@ 4.2 Results
OCRBench Results. We conduct a comparative analysis of our approach with recent large multimodal models. For our evaluation, we utilize three Scene Text-Centric VQA datasets: STVQA [51], TextVQA [33], and OCRVQA [52]; three Document-Oriented VQA datasets: DocVQA [36], InfoVQA [38], and ChartQA [37]; and three Key Information Extraction (KIE) datasets: FUNSD [53], SROIE [54], and POIE [55]. For a comprehensive assessment of performance, our evaluation includes OCRBench [12], a recent benchmark specifically developed to evaluate the Optical Character Recognition (OCR) capabilities of Large Multimodal Models. OCRBench spans a wide range of text-related visual tasks, encompassing 29 datasets, and is designed to generate an overall score.
Figure 5: Visualization results of TextMonkey. The bounding boxes generated by the model are visualized in red. The location of the ground truths are highlighted with green boxes.
@@
Table V: Quantitative accuracy of text spotting. The “Total-Text” and “CTW1500” datasets do not use a specific vocabulary for evaluation, while the “ICDAR 2015” dataset uses a general vocabulary for evaluation of other models. Note TTS only uses synthetic location data. TextMonkey is not fine-tuned by the downstream text spotting datasets without any vocabulary.
@@
TextMonkey 78.2 61.4 63.2 57.5 66.9 45.1
Document Benchmarks results. To further compare and assess the capabilities of our method, we conduct tests on additional datasets utilizing the specific evaluation metric provided in their paper: F1-score for Deepform and KLC, accuracy for WTQ, relaxed accuracy measure for ChartQA, ANLS for DocVQA, and VQA score for TextVQA.
The results, shown in Tab. IV, indicate that our model leads in performance on these datasets, outperforming other models. Across different domains, TextMonkey achieves a score of 71.5 in DocVQA, 30.6 in WTQ, 65.5 in ChartQA and 68.0 in TextVQA. It shows our model’s capability to handle documents, tables, charts, and scene text.
Text spotting results. To show the extensive capabilities of our model, we assessed its performance on text spotting datasets without finetuning, as detailed in Tab. V. Given our model’s focus on identifying complete text passages, we segmented the predicted content into individual words for analysis. We employed two evaluation methodologies to evaluate our model’s performance. In the “Trans” mode, text is considered correct if the answer contains this word. Conversely, the “Pos” mode requires the consideration of positional information in accordance with previous methods [62]. For both metrics, due to granularity issues of the output (TextMonkey often produces an integrated paragraph while others only produce desired words), the metric can not strictly follow the evaluation setup; however, both should be quite similar, as both the error and correct situations match in calculations.
To maintain TextMonkey’s consistent performance, we refrained from fine-tuning it with downstream text spotting data, unlike other methods that were optimized for either the “Trans” or “Pos” metrics. Our results reveal that, for the “Trans” metric, TextMonkey outperformed SPTS v2 by a margin of 10.9%. Regarding the “Pos” metric, it demonstrated competent text spotting capabilities, showing its ability in understanding both text content and spatial positioning.
4.3 Visualization
We conduct a qualitative evaluation of TextMonkey across various scenarios, including natural scenes and document images. As shown in the left part of Fig. 5, TextMonkey accurately locates and identifies text in both scene and document images. Besides, natural images in Fig. 5 (a), documents in Fig. 5 (b), charts in Fig. 5 (c), and tables in Fig. 5 (d) exemplify TextMonkey’s adeptness at discerning and interpreting visual and textual information within a wide range of scenarios. Overall, TextMonkey’s performance across diverse scenarios demonstrates its effectiveness in perceiving and comprehending textual information in various visual contexts.
@@ 4.4 Ablation Study
Ablation study on strategies of reducing token length. As demonstrated in Tab. VIII, substituting important tokens with random ones (without token filter) leads to an average decline in performance by roughly 12.7%. This decline is attributed to the increased complexity of optimizing random queries, which necessitates more datasets to achieve a generalized representation compared to utilizing significant tokens. Solely focusing on pivotal features (without resampler) and directly eliminating features incurs a loss of some information, showing a decrease in performance, such as a 2.1% drop in SROIE. Additionally, neglecting the order of tokens (with unsorted token filter) does not markedly impair performance, owing to the language model’s inherent ability to organize unordered tokens. Nevertheless, the lack of token order can still lead to decrease, especially evident in the result of DocVQA, with a 2.2% decrease in performance.
Interaction between input resolution and the number of tokens remained. As shown in Tab. IX, Directly increasing the resolution without compressing tokens can actually lead to consistent worse performance, especially with a decrease of 9.2% performance in DocVQA. We speculate that the increase in resolution results in a significant increase in redundant tokens, making it more difficult to find crucial information in our setting. Therefore, compressing tokens reasonably can lead to higher performance. Considering the sparsity of information in large-sized images, it is also necessary to consider selecting an appropriate value of “r” for different input resolutions. Besides, increasing the input resolution brings benefits to the dataset, which contains many large-sized images, with 0.2% performance gain for DocVQA and 3.2% performance gain for InfoVQA. However, for datasets like TextVQA and SROIE, which contain much smaller images, increasing the input resolution directly does not yield any gains.
Table IX: Interaction between resolution and the number of tokens remained “r”. “-” in “r” means do not use token resampler and keep all the remaining tokens.
@@
Figure 6: Examples of structuralization of chart and table using TextMonket†.
@@ 4.5 Structuralization
The structuralization of charts and tables holds substantial practical value. Structured charts and tables present data in a clear format, and by extracting structural information from images, computers can accurately parse and extract the data. This makes data analysis, statistics, and modeling more efficient and precise. It also helps reduce the complexity of information and improves its comprehensibility. As depicted in Fig. 6, our model is capable of structuring charts and tables into JSON format, demonstrating its potential for downstream applications. According to Tab. IV, TextMonkey† exhibits a performance improvement of 1.3% and 1.4% on tables and charts, respectively. This underscores that high-quality data not only enables the model’s structuralization capabilities but also amplifies the effectiveness of the related benchmarks. However, it is worth noting that this type of data will primarily benefit the data within its own domain, thus leading to a performance decrease for cross-domain TextVQA.
@@ 4.6 App Agent
Recently, there has been a lot of attention on using LMMs for the task of acting as agents for smartphone applications  [63, 64, 65]. Unlike existing intelligent phone assistants like Siri, which operate through system back-end access and function calls, this agent interacts with smartphone applications in a human-like manner, using low-level operations such as clicking and swiping on the graphical user interface (GUI). It eliminates the need for system back-end access, enhancing security and privacy as the agent does not require deep system integration. The GUI primarily consists of icons and text, and we explore the feasibility of TextMonkey on this aspect. We transformed 15k user click data from the Rico [66] dataset and performed downstream fine-tuning using TextMonkey. As qualitatively shown in Fig. 7, our model is able to understand user intent and click on the corresponding icons, which suggests the potential of the model to serve as an app agent by using downstream data.
@@
Figure 8: Examples of grounding the position of the answer.
5 Discussion
5.1 Interpretability
By examining the grounding information, we can identify the reasons behind the model’s errors, enhancing a better understanding of the model’s behavior. As shown in Fig. 8 (a), we ground to the white region, indicating that the model might be engaging in hallucination. We correctly identify the location but recognize it wrongly in Fig. 8 (b). Fig. 8 (c) highlights a scenario where the model grounds to incorrect text but still provides the correct answer. This could imply that there is some level of randomness or uncertainty in the model’s responses at this point. In Fig. 8 (d), the alignment between the position and text indicates that the model is more confident in its predictions. Therefore, based on these analyses, we can gain a better understanding of the model’s behavior and have a better awareness of the model’s hallucination, thus reducing the model’s hallucination.
5.2 Chain-of-Thought
We also conduct experiments on several datasets and observe inconsistent improvements if we require a model to provide the answer’s position, as shown in Tab. X. In datasets where the majority of answers are based on information within the images, such as DocVQA and SROIE, there is a noticeable benefit in requiring the model to provide the answer’s position. However, for datasets that involve reasoning tasks, such as ChartQA and InfoVQA, where questions require comparisons or quantitative analysis (e.g., "How much more is A than B?"), demanding positional answers can actually result in a detrimental effect. Upon further examination of the wrong answer, we consider that the requirement of grounding might have partially affected certain reasoning needs. Hence, it is essential to consider the nature of the dataset and the type of questions being asked when deciding whether to impose the requirement of positional answers.
Additionally, we believe that automating the process of constructing a thinking chain [67] in subsequent steps could be a promising direction for future research. By developing mechanisms to generate a coherent chain of reasoning automatically, we can potentially enhance the overall performance and reasoning capabilities of our models.
Table XI: Comparison with different shapes of bounding box.
Representation SROIE DocVQA TextVQA ChartVQA
Polygon 47.2 64.0 65.7 57.9
Rect 47.0 64.3 65.9 58.2
Point 47.9 65.0 66.0 58.3
5.3 Comparison Between Different Representations of Position
Recently, some methods [61] have used points to represent positions instead of rectangles and polygons. Firstly, intuitively, the cost of generating points during inference would be lower compared to generating rectangles and polygons, as generating Nx points is required for other forms of bounding boxes. We aim to further investigate and experimentally validate which form is more suitable for LMMs to learn. To maintain strict consistency in our experiments, we only applied transformations to the data while keeping the other training hyperparameters the same. For the points, we selected the center points of the bounding boxes that were the most meaningful.
As demonstrated in Table XI, employing points as visual cues significantly enhances performance over rectangles. In the case of Docvqa, there was an improvement of 0.7%, while for SROIE, the enhancement reached 0.9%. Furthermore, rectangles often surpass polygons in performance. This might be attributed to the previously discussed issue that redundant image tokens could increase the complexity of the model’s learning process. Similarly, extensive position representations might face comparable obstacles. Given these considerations, along with the associated inference costs, utilizing points as representations can be a viable strategy for appropriate tasks.
@@ 6 Conclusion
This paper introduces TextMonkey to address the challenges associated with text-heavy tasks such as document question answering and fine-grained text analysis. We adopt Shifted Window Attention with zero initialization to help establish relationships while increasing input resolutions using a sliding window. Increasing the resolution simultaneously increases the number of tokens. Through analyzing the redundancy of tokens, our proposed Token Resampler effectively reduces the number of tokens. Furthermore, by engaging in multiple text-oriented tasks simultaneously, TextMonkey enhances its perception and understanding of spatial relationships, leading to improved interpretability and support for clicking screen-shots. By comparing our model with various LMMs, our model achieved excellent results on multiple benchmarks. It is worth mentioning that we also find that directly increasing the input resolution does not always lead to improvements, particularly for much smaller images. This underscores the necessity of creating an efficient method for scaling resolution in documents where size changes can be dramatic.